OpenStudy (anonymous):

Part 1: Show the expanded form and simplification of x^6/x^2. Explain in your own words how you can simplify x^6/x^2 without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (allank):

Part 1: show the expanded form and simplification x^6/x^2 = xxxxxx/xx = xxxx Explain in your own words how to do it without writing the expanded form: "Subtract the exponents of the x^6 and x^2 to get x^(6-2) which is = x^4" Part 2: My own fraction: a^12/a^4 Simplification: a^(12-4) = a^8

OpenStudy (anonymous):

thanks!

OpenStudy (allank):

you're welcome :)